LHC 2011 - latest news

Week 15: Machine coordinators: Ralph Assmann and Bernhard Holzer

LHC coordination vistar




Link to report from morning meeting

Monday 18th April

  • 00:06 Beams dumped when injecting third 72b batch for beam 2. Kicker flash-over. Many BLM’s interlocked, LHCb BCM triggered, several magnets tripped, vacuum valves closed, cryo lost in 78 and 81.
  • Kicker had a spark after 1.5 micro-s, while injecting the 72b batch. First 36b part OK, wrong kick for second 36b part.
  • Since then: Recovery. Cryo OK. Vacuum OK. Collimators OK.

8:30 meeting

Sunday 17th April

  • 03:00 injection 336 bunches, ramp, squeeze luminosity.
  • 03:19 Stable beams #1718. Duration: 9h08.
    Peak lumi ATLAS/CMS: 3.8/3.6e32.
    Integrated lumi ATLAS/CMS/LHCb/ALICE: 10.1/9.6/4.3/0.06 pb-1.
  • 11:27 Lost QPS OK on RQD and RQF sector 3-4. Will reset at the end of next  rampdown.
  • 12:28 Beam dump. No orbit data for more than 10 min, SIS interlock.
  • 13:29 incorporating lumi scan : IP1 is working and IP5 in H only. IP5-V: request is out of range.
  • 13:39 Access started in CMS and US15 for ATLAS.
  • 14:58 CMS access completed.
  • 15:01 Electron stopper relay at Pt 4 not working => access to UJ43 not possible. Access piquet called.
  • 20:19 Electron stopper problem fixed (micro switch). Patrol lost. Pt 4 ( sectors 1,9,11), organize re-patrol within OP
  • 21:19 RQD QPS: MQ.33R3 reset by the MPE piquet.
  • 21:20 Vacuum valve fault on VVGSW.160.5L4.B fixed.
  • 23:35 Injection

9:00 meeting

Saturday 16th April

  • 00:20 Stable beams. 228b.
  • 05:15 Beam dump from electrical perturbation. In stable beams: 4h55
  • 08:00 Electrical perturbation in IR2 ongoing. IR2 compensator (active filter) to be disconnected.
  • 10:30 injection: 336 bunches, ramp, squeeze, lost due to orbit feedback
  • 13:00 injection 336 bunches, ramp, squeeze, instabilities, beam lost via BPM threshold
  • 16:00 injection 336 bunches, ramp, squeeze luminosity, dump: water failure
  • 17:39 Stable beams #1716. In stable beams: 1h40.
  • 19:19 Beam dump. Cut of demineralised water due to a faulty flow meter triggered PIC to interlock in S34.
  • injection 336 bunches, ramp, squeeze luminosity, dump: software interlock due to flaky BPM at primary collimator.

9:00 meeting

Friday 15th April

  • 00:00 machine setup for luminosity run
  • 01:58 stable run with 228 bunches
  • 11:00 luminosity leveling in IP8 tested successfully
  • 12:11 dump by EIC, precycle, injection & ramp. In stable beams: 10h13
  • 15:49 stable run 228 bunches
  • 19:18 Beam dump from power converter (PIC) RCBCH10.L4B2. In stable beams: 3h29
  • 20:45 Access for PC
  • 22:08 Injection

8:30 meeting

Thursday 14th April

  • 00h00: Interlock tests for ALFA and TOTEM without beam (SPS still down).
  • 01h33: SPS available again. Was a problem with radial steering function.
  • 03h58: Injection.
  • 04h55: Ramp.
  • 05h16: All OK from RF.
  • 05h26: Adjust.
  • 05h37: Beam losses during lumi-scan. IR6 BPM interlock (one bunch below threshold).
  • 06h12: ATLAS and CMS prepare ramp up. Delays in CMS.
  • 06h28: Sector 67 tripped due to glitch on 18kV.

8:30 meeting

Wednesday 13th April

  • Inject 48 bunches, ramp, squeeze. Luminosity optimisation, tune scan, rf scan to check robustness of machine settings.
  • 08h18: 228b fill dumped just before start of ramp. Reasons: Transverse damper switched off for chromaticity measurement.
  • 09h02: Injecting again for 228b. Beam dump due to losses at injection. SPS had slight reduction in scraping. PSB checks. Last bunches in batch had very low intensity. Still losses at 70% of dump threshold for 72b injection.
  • 13h12: Ramp.
  • Capture loss from DC-BCT with ramp 1704, 50ns, 228b. B1: loose 1.5E10 out of 2.8E13 (0.05%). B2: losse 2.0E10 out of 2.8E13 (0.07%).
  • Tune feedback switched off at 13:18:51 (due to a coupling that was larger than the tune split).
  • Losses associated with coupled bunch instability started at 13:19:00.
  • 13h33: Squeeze.
  • 14h57: Stable beams. Lumi 2e32 in ATLAS and CMS.
  • 16h37: Beams dumped by EIC. BPM studies start.
  • 17h46: Post-mortem. LBDS-B1 was armed during BPM calibration which triggered IR6 BPM interlock (no beam).
  • 18h17: SMP configuration changed to be able to obtain the STB and MDI flags at 450 GeV - for the TOTEM/ALFA tests tonight.
  • 18h21: Injection.
  • 19h29: Following the losses and observations done on the ramp this morning: Reduction of the chromaticity by 2 units for beam 1 and beam 2 - Horizontal plane in the middle of the ramp (see plots). Increased octupoles. Increased damper gain.
  • 20h19: BPM studies completed.
  • 20h46: RF problems in SPS.

8:30 meeting

Tuesday 12th April

  • 00h56: Access finished in pt 4, quench heaters are charged in cell 32R4.
  • 01h19: Access finishing in pt 3.
  • 02h40: Injection. 2+1b. Qualification of ramp cycle and new collimator functions:
  • 03h16: Ramp.
  • 03h37: Squeeze.
  • 03h58: Adjust
  • 04h46: Collisions optimized (except ALICE). Loss maps.
  • 06h11: Injection for scrubbing.

8:30 meeting

Monday 11th April

  • 07h34: beam with 1020b lost after 10 Hz RF trim. Fill 1694.
  • 07h56: XPOC to be reset by expert.
  • 08h18: Masked QPS_STATE SIS interlocks for RB.A12 and RB.A45.
  • 09h43: BPM calibration. BPM good and bad status flips.
  • 11h00: M. Andersen: “Did a quick fix of OFSU by changing Fesa threads priorities to allow the Orbit reading thread more CPU time. Back to the drawing board to determine if more substantial redesign is needed to make the system more robust.”
  • 11h25: LHC.USER.RAMP added to Hypercycle (RAMP_FAST_2011_V1 BP) for test ramp with 12+1 bunches + probe.
  • 11h42: XPOC problem.
  • 12h21: XPOC cleared. Start injection. Reduced ADT vertical gain of B1 & B2 by 50% (old value: 0.04 new value: 0.02). Reduced ADT gain for B1-V by another 50% to 0.01 (adapted 1st point of ADT - ramp functions as well).
  • 13h58: Start ramp. V.Baglin: All pressure interlock thresholds have been reset to their standard values (1e-7) for this test ramp.
  • 14h17: Flat top. Ramp was fine. Orbits at flat top recorded for bunch train (12b) and single nominal bunch. Squeeze and collision not possible due to required update of TCTH.4L2 function.
  • 15h26: Beams dumped.
  • 15h43: RCBH21.R4B2 tripped during ramp down. Reset.
  • 16h35: Sector 81 tripped. Bridge B cooling. Access. Give access for discharging quench heaters in S12 and S45. One QPS team only: do S12 first. CMS access in shadow.
  • 19h39: Access in tunnel point 4 for QPS – one of the electron stoppers does not move in, preventing QPS intervention. Calling experts. Cannot get in and out access anymore.
  • 20h57: Fire detection is faulty in point 3. Piquet is intervening.
  • 22h22: In shadow of access – TCT collimator functions for squeeze and physics beam processes updated. The nominal sequences are fully operational.
  • 22h29: Access team intervention in IR4 completed. Relais was broken.

8:30 meeting